In a survey, 85% of the voters support a particular referendum. If 40 voters are chosen at random, and X is the number of voters that support this referendum, find the standard deviation X.
Draw a graph showing the probability distribution for the sample space for tossing four coins.
On a Saturday evening, 34% of the people in Chicago go out to dinner, 18% see a movie, 13% have a party, and 35% stay home. Seventeen people are randomly selected. Can the probability that exactly 4 of them stay home be computed using binomial model?
Please help with questions. Need ASAP. Thank you in advance.

Answer by stanbon(75887)

In a survey, 85% of the voters support a particular referendum.
If 40 voters are chosen at random, and X is the number of voters
that support this referendum, find the standard deviation X.
-------
std = sqrt(npq) = sqrt(40*0.85*0.15) = 2.26
----------------------------------------------------
Draw a graph showing the probability distribution for the sample
space for tossing four coins.
4 heads::: 4C4/16 = 1/16
3 heads::: 4C3/16 = 4/16
2 heads::: 4C2/16 = 6/16
1 head:::: 4C1/16 = 4/16
0 heads::: 4C0/16 = 1/16
=========================
On a Saturday evening, 34% of the people in Chicago go out to dinner, 18% see a movie, 13% have a party, and 35% stay home. Seventeen people are randomly selected. Can the probability that exactly 4 of them stay home be computed using binomial model?
-----
Binomial?? No
The probability for each of the 17 is not the same.
====================================================
Cheers,
Stan H.